x-2y=2
2x+3y=11
use substitution method to solve system
if dependant or inconsistant please indicate

Answer by fractalier(6550)

From
x - 2 y = 2
2x + 3y = 11
Solve for x in the first and substitute into the second:
x = 2y + 2
2(2y + 2) + 3y = 11
4y + 4 + 3y = 11
7y + 4 = 11
7y = 7
y = 1
Then since x = 2y + 2, we have
x = 2(1) + 2
x = 4
Our solution is (4, 1)